OpenStudy (anonymous):

How do you solve this polynomial? 2(x^3)-x^2-1=0 I know how to solve polynomials, but I can't remember how to solve when you can't factor normally.

OpenStudy (blurbendy):

you can simplify this to: (x-1)(2x^2 + x + 1) = 0

OpenStudy (blurbendy):

your solution will depend on if your looking for just the real solutions, or the complex solutions as well

OpenStudy (anonymous):

Could you explain to me how you got there? I'm sorry, I just want to learn how to do it too XD

OpenStudy (anonymous):

And does that say two x squared DIVIDED by x DIVIDED by 1?

OpenStudy (anonymous):

x=1 is a root. then x=1 is a factor. and then factorise.

OpenStudy (blurbendy):

quantity (x - 1) times (2x^2 + x + 1) = 0

OpenStudy (blurbendy):

no division

OpenStudy (anonymous):

Thanks! But how did you find out that x=1 was a factor?

OpenStudy (blurbendy):

(x-1)(2x + x + 1) = 0 consider first (x - 1) = 0 x = 1

OpenStudy (blurbendy):

that's your real solution. you'll get complex solutions if you try to find 2x^2 + x + 1 = 0

OpenStudy (anonymous):

Okie dokie. Thanks so much for your help!

OpenStudy (blurbendy):

yw. hope it makes sense